A psychologist wishes to estimate the true proportion of all Americans who dislike going to the dentist.

How many Americans should be surveyed in order to be 98 percent confident that the sample mean will not differ from the population mean by more than 3% (.03)
Answer #1

Solution: Given that level of confidence = 9870 = 0.98 L=1-098 = 10.02 42 = 0.02/2 - 10.01 Zd/2 = 2.3 0.03 population proport